# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ Abstract
The “free trial” followed by automatic renewal is a dominant business model in the digital economy. Standard models explain trials as a mechanism for consumers to learn their valuation for a product. We propose a complementary theory based on rational inattention. Consumers know their valuation but face a cognitive cost to remember to cancel an unwanted subscription. We model this using a Shannon entropy-based cost of information processing, where a consumer’s baseline attention level decays with the length of the trial period. This creates a novel trade-off for a monopolist firm: a longer trial increases “inattentive revenue” from consumers who fail to cancel, but it also lowers ex-ante consumer utility, making the initial offer less attractive. We show that this trade-off leads to an interior optimal trial length, even for products where value-learning is instantaneous. Our model, under standard assumptions about demand elasticity and the distribution of consumer valuations, generates sharp, testable predictions about the relationship between contract terms. We find that the optimal renewal price and trial length are complements: firms offering longer trials will also set higher post-trial prices. We analyze the impact of policies aimed at curbing consumer exploitation, such as “click-to-cancel” regulations. We show that such policies, by making attention effectively cheaper, lead firms to reduce trial lengths. The effect on price depends directly on the elasticity of demand from loyal subscribers. We also extend the model to include paid trials, showing that introductory prices and trial lengths act as strategic substitutes. Our framework provides a micro-founded explanation for common features of subscription contracts and offers a new lens through which to evaluate consumer protection policies in digital markets.
Keywords: Rational Inattention; Free Trials; Subscription Contracts; Digital Services; Behavioral IO.
JEL Codes: D11, D83, D91, L12, M31.
@@ 1 Introduction
The proliferation of subscription-based services has become a defining feature of the modern economy. From streaming entertainment (Netflix, Spotify) and software as a service (Adobe Creative Cloud, Microsoft 365) to news media and e-commerce (Amazon Prime), the automatic renewal model is ubiquitous. A central approach in the customer acquisition and growth strategy toolkit for these services is an introductory trial offer, typically a free or low-cost period, after which the consumer is automatically enrolled in a recurring payment plan (commonly referred to as the “negative option” model by regulatory bodies). The subscription mode accounts for nearly half the share of mobile apps revenue, despite the small share of subscription-based apps. The global subscription e-commerce market has grown into a multi-billion-dollar industry, underscoring the strategic importance of understanding the design of these contracts.
@@
The dominant theoretical explanation for trial offers in economics and marketing is rooted in the problem of experience goods (Villas-Boas, 2006). In this view, consumers are uncertain about their valuation for a new product. A free trial serves as a mechanism to resolve this uncertainty, allowing consumers to learn their personal fit and make an informed purchase decision. The firm offers the trial to signal quality and persuade high-valuation consumers to adopt, balancing the cost of the trial against the future stream of revenue from converted subscribers. The optimal length of such a trial is thus determined by the time required for this learning process to conclude.
While the learning motive is undeniably important, it provides an incomplete picture of the subscription landscape. We observe trial periods that seem excessively long relative to any plausible learning period. A consumer can likely assess a music streaming service’s library and interface within a few days, yet 30-day, 60-day, or even longer trials are standard practice. Furthermore, a substantial body of evidence points to consumer inertia, forgetfulness, and procrastination as significant drivers of subscription retention. Surveys consistently reveal that a large number of consumers pay for subscriptions they no longer use, with forgetfulness cited as a primary reason. This suggests that trial periods may serve a dual purpose: one of facilitating learning, and another, perhaps less benign, of leveraging consumer inattention.
@@
This paper develops a formal theory of subscription contract design that isolates and explores this second channel. We propose a model where the trial length’s primary strategic function is to modulate consumer attention. To achieve a clean separation from the learning motive, we assume consumers know their valuation for the service from the outset. The central friction in our model is not uncertainty about value, but the cognitive cost of future actions. A consumer who signs up for a trial and finds the service is not worth the renewal price must remember to perform a specific action, cancellation, by a specific deadline. This act of timely recall is not costless; it requires cognitive resources.
@@
We formalize this friction using the rational inattention framework pioneered by Sims (2003) and adapted for discrete choice by Matějka and McKay (2015). In our model, a consumer who wishes to cancel can exert mental effort to increase the probability of successfully doing so. The cost of this effort is governed by their "attention sensitivity." Our key behavioral assumption, grounded in the psychology of prospective memory, is that this attention sensitivity decays with the length of the trial period. A task due tomorrow is more salient and easier to manage than a task due three months from now.
@@
This simple and intuitive assumption generates a novel and fundamental trade-off for a profit-maximizing firm. On one hand, extending the trial period causes consumer attention to decay. This makes it more costly for low-valuation consumers to ensure they cancel, leading to a higher rate of "accidental" renewals. This generates what we term inattentive revenue, a stream of profit derived purely from consumer forgetfulness. This creates a powerful incentive for the firm to lengthen the trial period, an incentive entirely absent from standard learning models.
@@
On the other hand, the consumers in our model are fully rational and sophisticated. They are aware of their own cognitive limitations and anticipate the challenges of managing a subscription with a distant deadline. They understand that a longer trial increases the expected cost of monitoring and the risk of a costly mistake. This prospective cost lowers their ex-ante expected utility from signing up for the contract in the first place. We label this effect the IR-Slack. It acts as a countervailing force, making the firm’s offer less attractive and providing a strong incentive to shorten the trial period.
The optimal contract, comprising a trial length ($T$) and a renewal price ($P$), is designed to resolve this tension. Our first major result demonstrates that this trade-off naturally gives rise to a finite, interior optimal trial length. The model thus explains why firms offer trials that are longer than necessary for learning but are not infinitely long.
@@
We then deploy this framework to generate sharp, testable predictions about the structure of subscription contracts. By imposing standard, empirically-grounded assumptions on the distribution of consumer valuations, we find that trial length and renewal price are strategic complements. A firm that lengthens its trial period to enhance its inattentive revenue stream will also find it optimal to increase its renewal price to extract more value from this channel. This prediction provides a clear point of departure from learning-based models, where the relationship between trial duration and price is ambiguous.
@@
Our framework also serves as a powerful tool for analyzing public policy. Growing regulatory concern about "dark patterns" and exploitative business practices has led to proposals like "click-to-cancel" laws. We model such a policy as an exogenous shock that makes attention effectively cheaper for all consumers. We show that this forces firms to compete less on the inattention margin, leading to shorter optimal trial lengths. The effect on price, however, depends directly on the price elasticity of demand from the firm’s base of willing subscribers. This result highlights the potential for well-intentioned regulations to have unintended consequences, underscoring the need for careful, theory-informed policy design.
@@
Finally, we extend our baseline model to consider paid trials. We show that the introductory price and the trial length act as strategic substitutes. An introductory price functions as a screening device, selecting for more committed or higher-valuation consumers. When a firm uses this tool, it diminishes the pool of low-valuation consumers who are the source of inattentive revenue, thereby diluting the incentive to use a long trial. This "dilution principle" illustrates how different contract features can substitute for one another in a firm’s strategy to manage consumer behavior.
@@
This paper contributes to three distinct streams of literature. First, we contribute to the rational inattention literature by endogenizing the cost of information; the firm’s strategic choice of trial length determines the consumer’s cognitive environment. Second, we advance the behavioral industrial organization literature on exploitation contracts by modeling a market with sophisticated consumers who are aware of their cognitive frictions, providing a new micro-foundation for consumer inertia. Third, we offer a new, complementary theory for the economics of subscriptions and marketing strategy, focusing on attention rather than learning as a primary driver of contract design.
@@
The paper proceeds as follows. Section 2 provides a detailed review of the related literature. Section 3 formally lays out the model of the consumer’s and firm’s problems. Section 4 presents our main results that hold for a general class of information cost functions. Section 5 specializes the model to the widely used Shannon entropy cost structure and imposes regularity conditions to derive sharper predictions. Section 6 analyzes the extension to paid trials. Section 7 synthesizes the findings into a discussion of managerial and policy implications. Section 8 concludes and outlines avenues for future research.
@@ 2 Related Literature
This research is positioned at the intersection of three influential streams of literature: the economic theory of rational inattention, behavioral industrial organization with a focus on contract design, and the marketing and information systems (IS) literature on digital subscriptions and customer management. We synthesize insights from each to build our theory of contract design based on attentional friction.
@@ 2.1 Rational Inattention and Bounded Rationality in Choice
The theory of rational inattention, introduced by Nobel laureate Sims (1998, 2003), provides a rigorous micro-foundation for information frictions. It posits that agents are not perfectly informed, not because information is unavailable, but because processing it is cognitively costly. Agents must therefore optimally allocate their finite attention. Our work uses the tractable discrete-choice formulation of Matějka and McKay (2015), where Shannon entropy models the cost of distinguishing between alternatives, leading to a generalized logit choice model.
@@
While much of the early work in rational inattention focused on macroeconomics, its principles are increasingly applied to microeconomic and marketing contexts. For instance, it provides a theoretical underpinning for why consumers might not react to small price changes or may exhibit inertia in their choices (Maćkowiak and Wiederholt, 2009). Our primary contribution to this literature is to endogenize the cost of attention. In most RI models, the agent’s information processing capacity (our $\lambda_{0}$) is an exogenous parameter. In our model, the effective attention sensitivity $\lambda(T)$ is a direct function of a firm’s strategic choice: the trial length $T$. This connects our work to a growing interest in how market environments and firm strategies can shape consumer cognition. For example, de Clippel et al. (2014) study how firms might design complex products to exploit consumers with limited attention. Our model provides a new mechanism, temporal distance, through which firms can strategically influence the cognitive burden on consumers. By making the consumer’s attention capacity an outcome of the firm’s contract design, we create a feedback loop between market structure and cognitive constraints, allowing us to study how firms might strategically manipulate the informational environment to their advantage.
2.2 Behavioral IO and Exploitation Contracts
@@
Our paper is situated within the field of behavioral industrial organization, which examines market outcomes when firms interact with consumers who exhibit psychological biases. A significant strand of this literature focuses on the design of "exploitation contracts" that generate profit from consumer mistakes. A seminal contribution is Gabaix and Laibson (2006), who study markets with shrouded attributes. They show that firms may choose to hide the price of add-ons (e.g., printer ink) to exploit "myopic" consumers. Our model shares the theme of firms profiting from a secondary, often overlooked, aspect of a product, in our case, the cost of cancellation.
@@
However, our conception of the consumer is different and aligns more closely with recent behavioral models that assume consumer sophistication. Rather than positing a fraction of naive consumers, all consumers in our model are sophisticated and forward-looking. They are "rationally inattentive," not naively forgetful. They are fully aware of their own cognitive limitations and anticipate that a long trial will make cancellation more difficult. This anticipation is precisely what generates the "IR-Slack," the force that disciplines the firm’s incentive to exploit inattention.
@@
In this respect, our work is closer in spirit to models of contracting with sophisticated but time-inconsistent agents, such as those with present-biased preferences (DellaVigna and Malmendier, 2004; Heidhues and Kőszegi, 2017). DellaVigna and Malmendier (2004) analyze contracts for gyms and credit cards, showing that firms will optimally design contracts to take advantage of consumers’ self-control problems, even when consumers are sophisticated. Our paper provides a parallel analysis where the friction is inattention rather than present bias. We thus contribute to this literature by proposing a new, attention-based mechanism for consumer inertia and showing how it can be embedded within a contract design problem with sophisticated agents. This approach also resonates with the marketing literature on "sludge," which examines how firms can increase cognitive friction to guide consumer choices, often to the consumer’s detriment (Thaler and Sunstein, 2008).
@@
Our theoretical framework finds strong empirical support in a recent large-scale field experiment by Miller et al. (2023). In a study with over two million newspaper readers, they experimentally vary the presence of an automatic renewal clause. Their findings validate the two central, opposing forces in our model. First, they document powerful inertia, finding that a large fraction of consumers who are defaulted into a paid subscription remain subscribed while exhibiting little to no usage, which corresponds to the inattentive revenue channel we model. Second, and critically, they find strong evidence of consumer sophistication, showing that initial take-up of auto-renewal offers is 24%–36% lower than for equivalent auto-cancel offers. This provides a direct empirical measurement of the choice-based deterrence that we label the IR-Slack. Their estimate that a majority of inert consumers are sophisticated and anticipate their future inertia lends crucial support to our decision to model consumers as rational agents who are aware of their own cognitive frictions, rather than as naively myopic agents.
2.3 Marketing Strategy and Subscription Models
There is a vast literature in marketing and related business fields on pricing and promotion strategies for digital goods. The classic explanation for free trials is to facilitate consumer learning and reduce perceived risk for experience goods (Villas-Boas, 2006). In this context, trials are a form of promotional spending to drive adoption. For instance, Narayan et al. (2011) study how seeding strategies and free trials can be used to launch new products in the presence of network effects. Our model provides a complementary, attention-based rationale that can explain trial lengths that extend beyond any plausible learning period. These two mechanisms, learning and attention, are not mutually exclusive. A complete model of subscription trials would likely include both, but our contribution is to isolate the attention channel and derive its unique strategic implications.

### v3

# Automatic Renewal Trials and Rationally Inattentive Customer

@@ Abstract
The “free trial” followed by automatic renewal is a prevalent business model in the digital economy. While traditional models view trials as a way for consumers to learn a product’s value, we propose a complementary theory based on rational inattention (Sims, 2003). In our framework, consumers already know their valuation but face cognitive costs associated with remembering to cancel an unwanted subscription. We model this using a Shannon attention cost, where a consumer’s attention decays over the trial period. This creates a fundamental trade-off for firms: longer trials increase “inattentive revenue” from consumers who forget to cancel, but they also reduce the initial attractiveness of the offer. This dynamic leads to an optimal trial length, even when consumers do not need time to learn the product’s value. Our model predicts that optimal renewal prices and trial lengths are complements; longer trials are associated with higher post-trial prices. Analyzing consumer protection measures, such as click-to-cancel regulations, we find that making cancellation easier motivates firms to shorten trial periods. Furthermore, we demonstrate that introductory prices (paid trials) and trial lengths act as strategic substitutes. This framework provides a new lens for understanding subscription contracts and evaluating consumer protection policies in digital markets.
Keywords: Rational Inattention; Free Trials; Subscription Contracts; Digital Services; Pricing.
@@ 1 Introduction
“The struggle of man against power is the struggle of memory against forgetting.”
Milan Kundera
The proliferation of subscription-based services has become a defining feature of the modern economy. From streaming entertainment (e.g., Netflix, Spotify) and software as a service (Adobe Creative Cloud, Microsoft 365) to news media and e-commerce (Amazon Prime), or even meal preps (e.g., Home Chef, Factor), the “free trial with automatic paid subscription renewal” model is ubiquitous. A central approach in the customer acquisition and growth strategy toolkit for these services is an introductory trial offer, typically a free or low-cost period, after which the consumer is automatically enrolled in a recurring payment plan (commonly referred to as the “negative option” model by regulatory bodies). The subscription mode accounts for almost half the share of mobile apps revenue, despite the small share of subscription-based apps. Given the multi-billion-dollar scale of this market, understanding the strategic design of these contracts is highly valuable.
The dominant theoretical framework to examine product trial offers in existing literature follows the paradigm of consumer learning (Hoch and Ha, 1986; Iyengar et al., 2007). In this sense, consumers are uncertain about their valuation of a new product. A free trial serves as a mechanism to resolve this uncertainty, allowing consumers to learn their personal fit and make an informed purchase decision. The firm offers the trial to signal quality and persuade high-valuation consumers to adopt, balancing the cost of the trial against the future stream of revenue from converted subscribers. The optimal duration of such a trial is determined by the time required for this learning process to conclude.
While consumer learning is undeniably an important factor, it provides an incomplete picture of the subscription landscape. We often observe trial periods that appear excessively long, relative to any reasonable learning period. A user can likely assess a music streaming service’s library and interface within a few days, yet 30-day, 60-day, or even longer trials are standard practice. Furthermore, a substantial body of evidence points to consumer inertia, forgetfulness, and procrastination as significant drivers of subscription retention. Surveys consistently reveal that a large number of consumers pay for subscriptions they no longer use, with forgetfulness cited as a primary reason (nearly half, according to a recent Forbes survey Orentas and Allen 2024). This suggests that trial periods may serve a dual purpose: one of facilitating learning, and another, perhaps less benign, of exploiting consumer inattention.
@@
We develop a formal theory of subscription contract design, in the case of auto-renewal, that isolates and explores this second channel. We propose a model where the primary strategic function of the trial length is to modulate consumer attention. To separate our mechanism from consumer learning, we assume that consumers know their valuation of the service early on [^1] As discussed later in the paper, these two factors are likely complementary, and our assumption is purely to keep the model parsimonious.. The central friction in our model is not uncertainty about value, but the cognitive cost of future actions. A consumer who signs up for a trial and finds that the service is not worth the renewal price must remember to perform a specific action, cancellation, by a specific deadline. This act of timely recall is not costless; it requires cognitive resources.
@@
We formalize this friction using the rational inattention framework pioneered by Sims (2003) and adapted for discrete choice by Matějka and McKay (2015). In our model, a consumer who wishes to cancel can exert mental effort to increase the probability of successfully doing so. The cost of this effort is governed by their "attention sensitivity." Our key behavioral assumption, grounded in the psychology of prospective memory (Krishnan and Shapiro, 1999; McDaniel and Einstein, 2000), is that this attention sensitivity decays with the length of the trial period. A task due tomorrow is more salient and easier to manage than a task due three months from now.
@@
This intuitive assumption generates a novel and fundamental trade-off for the firm. On one hand, extending the trial period encourages attentional decay. This makes it cognitively more expensive for low-valuation consumers to ensure cancellation, leading to a higher rate of accidental renewals. This generates what we term inattentive revenue, i.e., profit derived purely from consumer forgetfulness. This creates a strong incentive to lengthen the trial, a mechanism absent in standard learning models. On the other hand, the consumers in our model are sophisticated. They rationally anticipate their own cognitive limits and the associated costs of monitoring a distant deadline. A longer trial increases the risk of a costly mistake, lowering their ex-ante expected utility from signing up. We term this deterrent effect IR-Slack. It acts as a powerful balancing force, incentivizing the firm to shorten the trial to increase initial adoption.
@@
On the other hand, the consumers in our model are fully rational and sophisticated. They are aware of their own cognitive limitations and anticipate the challenges of managing a subscription with a distant deadline. They understand that a longer trial increases the expected cost of monitoring and the risk of a costly mistake. This potential cognitive burden lowers their expected utility from signing up for the subscription in the first place. We label this component the IR-Slack, and this acts as a constraint, making the firm’s offer appear less attractive and providing a strong incentive to shorten the trial period.
@@
The optimal contract, consisting of the trial length ($T$) and renewal price ($P$), balances this tension, resulting in moderate-length trials we often observe in practice. The model thus explains why firms offer trials that are longer than necessary for learning but not exceedingly so. We then employ this framework to generate testable predictions about the structure of subscription contracts. By imposing standard assumptions on the distribution of consumer valuations, we find that trial length and renewal price are strategic complements. A firm that lengthens its trial period to enhance its inattentive revenue stream will also find it optimal to increase its renewal price to extract more value from this channel. This prediction diverges from the consumer learning framework, where the relationship between trial length and price is often unclear.
@@
Our model also provides a lens for analyzing regulatory interventions. Concerns over “dark patterns” have spurred regulator proposals like “click-to-cancel” laws, recently introduced in California and proposed federally by the FTC (Luguri and Strahilevitz, 2021). We model such policies as an exogenous reduction in the cost of consumer attention. We show that this forces firms to retreat from the inattention margin, leading to shorter optimal trial lengths. The impact on pricing, however, is nuanced, depending crucially on the price elasticity of demand among willing subscribers. This highlights the necessity of theory-informed policy design to anticipate unintended consequences.
@@
Finally, we extend our baseline model to the case of paid trials. We show that the introductory price and the trial length act as strategic substitutes. An introductory price functions as a screening device to select for more committed, potentially higher CLV customers. This plan design will reduce the pool of “low-valuation” consumers who are the source of inattentive revenue, therefore diluting the incentive to use a long trial. This “dilution principle” illustrates how different contract features can substitute for each other in a firm’s strategy to manage consumer behavior.
@@
Our study contributes to three distinct strands of literature. First, we advance the rational inattention literature by endogenizing the cost of attention, wherein the firm’s contract design actively shapes the consumer’s cognitive environment. Second, we contribute to behavioral industrial organization by providing a new micro-foundation for consumer inertia in a market populated by sophisticated consumers. Third, we offer a novel framework for business model strategy, positioning attention, rather than learning, as a central driver of subscription contract design.
@@
The rest of the paper proceeds as follows. First, Section˜2 provides a detailed review of the related literature. Section˜3 formally introduces the primitives of the model, as well as the description of the consumer’s and firm’s problems. In Section˜4 we present our main results that hold for a general class of information cost functions. Next, Section˜5 specializes the model to the widely used Shannon entropy cost structure and imposes regularity conditions to derive sharper predictions. Additionally, in Section˜6 we extend the model to the (less but still common) paid trials setting. In Section˜7, we synthesize the findings into a discussion of managerial and policy implications. Finally, Section˜8 concludes and outlines avenues for future research. Detailed proofs of our claims are provided in the accompanying Appendices.
@@ 2 Related Literature
This research sits at the intersection of three influential streams of literature: the economic theory of rational inattention, behavioral industrial organization with a focus on contract design, and the literature on digital subscriptions and customer management. By synthesizing insights from each domain, we build a comprehensive theory of contract design based on attentional friction that bridges economic theory with managerially relevant practice.
@@ 2.1 Rational Inattention and Bounded Rationality in Choice
The foundation of our theoretical framework rests on the theory of rational inattention, introduced by Sims (1998, 2003). This theory provides a rigorous micro-foundation for information frictions, positing that agents are not perfectly informed, not because information is unavailable, but because processing such information is cognitively costly. Agents must therefore optimally allocate their finite attention. Our work builds on the tractable discrete-choice formulation of Matějka and McKay (2015), where Shannon entropy models the cost of distinguishing between alternatives, leading to a generalized logit choice model.
While early applications of rational inattention focused primarily on macroeconomic phenomena, its principles have increasingly found application in microeconomic and business contexts. The framework provides theoretical underpinning for consumer behaviors such as non-responsiveness to small price changes and choice inertia (Maćkowiak and Wiederholt, 2009). Several studies have applied rational inattention to explain strategic pricing behavior, including rigid pricing when consumers are inattentive to price (Matějka, 2015), discrete pricing when sellers are inattentive (Matějka, 2016), and quality-based strategies when consumers are inattentive to quality (Martin, 2017). Our contribution to this stream of research is examining the dynamic case when consumers are inattentive to timing, a novel yet extremely common and applicable setting in subscription markets. This temporal dimension of inattention has received limited attention in prior work, despite its clear relevance to modern digital commerce.
Recent methodological advances have made rational inattention models more empirically tractable. Joo (2023) develop a new framework for empirical discrete choice demand estimation based on Matějka and McKay (2015)’s work, while Brown and Jeon (2024) introduce a closed-form tractable model of RI-based discrete choice through the elegant assumption of Cardell distribution shocks. Readers interested in empirical applications in marketing and related fields should refer to Turlo et al. (2025) for a detailed overview. These developments provide the empirical foundation for testing our theoretical predictions. Empirical researchers can easily extend these existing approaches to our framework by incorporating a dynamic state-space component.
@@
A major theoretical contribution of our paper is endogenizing the cost of attention. In most extant RI models, the agent’s information processing capacity (our $\tau_{0}$) is treated as an exogenous parameter. In our model, the effective attention sensitivity $\tau(T)$ becomes a direct function of a firm’s strategic choice: the trial length $T$. This innovation connects our work to growing interest in how market environments and firm strategies can shape consumer cognition (de Clippel et al., 2014). By making the consumer’s attention capacity an outcome of the firm’s contract design, we create a feedback loop between market structure and cognitive constraints, enabling analysis of how firms might strategically manipulate the informational environment to their advantage.
2.2 Behavioral Industrial Organization and Exploitation Contracts
@@
Our paper contributes to the rapidly growing field of behavioral industrial organization, which examines market outcomes when firms interact with consumers who exhibit psychological biases. A significant strand of this literature focuses on the design of "exploitation contracts" that generate profit from consumer mistakes. The seminal work of Gabaix and Laibson (2006) on markets with shrouded attributes demonstrates how firms may hide the price of add-ons (e.g., printer ink) to exploit "myopic" consumers. Our model shares the theme of firms profiting from a secondary, often overlooked aspect of a product, which in our case is the cognitive cost of cancellation.
@@
However, our conception of the consumer differs fundamentally from models that assume consumer naivety. Rather than positing a fraction of naive consumers, all consumers in our model are sophisticated and forward-looking. They are "rationally inattentive," not naively forgetful. Consumers are fully aware of their own cognitive limitations and anticipate that a long trial will make cancellation more difficult. This anticipation generates what we term "IR-Slack": a counterbalancing loss component that moderates the firm’s incentive to fully exploit inattention.
@@
This approach aligns more closely with recent behavioral models that assume consumer sophistication, particularly those dealing with time-inconsistent agents such as those with present-biased preferences (DellaVigna and Malmendier, 2004; Heidhues and Kőszegi, 2017). DellaVigna and Malmendier (2004) analyze contracts for gyms and credit cards, demonstrating that firms will optimally design contracts to exploit consumers’ self-control problems, even when consumers are sophisticated. Our paper provides a parallel analysis where the friction is inattention rather than present bias, contributing to this literature by proposing a new, attention-based mechanism for consumer inertia and showing how it can be embedded within a contract design problem with sophisticated agents. Recent empirical evidence by Rodemeier (2025) lends credence to our framework, as that study shows, through a field experiment, that consumers are sophisticated about their inattention in rebate redemption.
This approach also resonates with the behavioral economic literature on "sludge," which examines how firms can increase cognitive friction to guide consumer choices, often to the consumer’s detriment (Thaler and Sunstein, 2008). Our model formalizes these intuitions within a rigorous economic framework, providing testable predictions about when and how firms will deploy such strategies.
@@
Empirical validation of our theoretical framework comes from a recent large-scale field experiment by Miller et al. (2023). In a study with over two million newspaper readers, they experimentally vary the presence of an automatic renewal clause. Their findings validate the two central, opposing forces in our model. First, they document powerful inertia, finding that a large fraction of consumers who are defaulted into a paid subscription remain subscribed while exhibiting little to no usage, corresponding to the inattentive revenue channel we model. Second, and critically, they find strong evidence of consumer sophistication, showing that initial take-up of auto-renewal offers is 24%–36% lower than for equivalent auto-cancel offers. This provides direct empirical measurement of the choice-based deterrence that we label the IR-Slack. Their estimate that a majority of inert consumers are sophisticated and anticipate their future inertia lends crucial support to our decision to model consumers as rational agents aware of their own cognitive frictions.
2.3 Digital Economy, Subscription Models, and Customer Management
The literature on digital economy and subscription models provides a crucial context for our theoretical framework. The traditional explanation for free trials centers on facilitating consumer learning and reducing perceived risk for experience goods (Israel, 2005; Iyengar et al., 2007). In this sense, trials represent a form of promotional spending to drive adoption. Lee and O’Connor (2003) demonstrate how seeding strategies and free trials can launch new products in the presence of network effects, while Wang and Özkan-Seely (2018) show how free trials can signal quality in markets with asymmetric information.
Our model provides a complementary rationale based on Rational Inattention that can explain trial lengths that extend beyond any plausible learning period. The learning and attention mechanisms are not mutually exclusive, and in fact, a complete model of subscription trials would likely include both. However, our contribution is to isolate the attention channel and derive its unique strategic implications, particularly relevant in markets where product quality is easily observable but cancellation requires active effort. As far as we know, ours is the first paper to establish a theoretical connection between subscription model and rational inattention, and the empirical literature on this connection is still highly sparse, with only one recent field experiment by Einav et al. (2025).
The freemium model literature offers additional insight into our setting. For example, Pauwels and Weiss (2008) and Cao et al. (2023) analyze drivers of conversion from free to paid tiers, while several other studies such as Lee et al. (2017) and Li et al. (2019) examine how firms can optimize the freemium experience to maximize long-term value, by balancing the trade-off between growth and monetization. Our focus on trial length as a strategic variable complements these studies by highlighting an underexplored dimension of freemium strategy. Similarly, research on digital platform engagement (Yoganarasimhan et al., 2023) and user stickiness (Ray et al., 2012; Zhou et al., 2024) provides behavioral foundations for understanding why consumers might remain subscribed despite low usage.